OpenStudy (anonymous):

A clock radio is rated as 30 W of power output. If the radio also draws 30 W at 120 V, which will the current draw be?

OpenStudy (anonymous):

how do you find it, anyone???

OpenStudy (anonymous):

\[\huge P=V \times I\] Where, P is power V is voltage I is current

OpenStudy (anonymous):

so it'll be 4A??

OpenStudy (anonymous):

No, try again.

OpenStudy (anonymous):

speed = 100 million m/s and frequency = 50 million Hz speed = 150 million m/s and frequency = 1500 Hz speed = 300 million m/s and frequency = 100 Hz what about this one??

OpenStudy (anonymous):

ok ill try again

OpenStudy (anonymous):

Once you got it, let me know. Then close this thread and post your question in a new thread.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im not sure im stuck on it

OpenStudy (anonymous):

30 = 120 x I Solve for I